What are the dimension of the first rectangle? if in a certain rectangle,the sum of the length and the width is 14 (measured in cm). A second rectangle is 5 times as look as the first, and 3 times as wide. The perimeter of the second rectangle is 91
Answer by stanbon(75887)

What are the dimensions of the first rectangle? if in a certain rectangle,the sum of the length and the width is 14 (measured in cm).
length = x
width = 14-x
-----------------------------
A second rectangle is 5 times as long as the first, and 3 times as wide.
length = 5x
width = 3(14-x) = 42-3x
---------------------------------
The perimeter of the second rectangle is 91
2[5x + 42-3x] = 91
------
2[2x+42] = 91
----
4x + 84 = 91
4x = 7
x = 7/4 (length of the 1st rectangle)
14-x = 14-(7/4) = 12 1/4 (width of the 1st rectangle)
=================
Cheers,
Stan H.
------------------
----